OpenStudy (tylermckinney16):

You plan to rent a car for your road trip. If you will be gone for 5 days and plan to drive 650 miles, which of these is the better deal? Explain your answer. Company A: $20 per day; unlimited mileage Company B: $15 per day; $0.20 per mile driven

OpenStudy (tylermckinney16):

@jdoe0001

OpenStudy (mathmale):

You need to calculate the TOTAL cost of renting the car for 5 days, 650 miles. Each rental company charges you a daily rate, PLUS a per-mile rate. You need to develop an equation for each of the 2 companies. It must be in dollars.

OpenStudy (mathmale):

Note that one of the 2 companies does NOT charge a per-mile fee. You can still develop a separate equation for both companies. Substitute 5 for the number of days and 650 for the total number of miles driven.

OpenStudy (mathmale):

Company A: $20 per day; unlimited mileage. this is the company that does not charge a per-mile fee. How much would Company A charge for a 5-day rental?

OpenStudy (jdoe0001):

well, let's see both plans \(\bf \begin{array}{llllll} &days&mileage&cost \\\hline\\ \textit{plan A}&{\color{brown}{ 5}}& {\color{blue}{ 650}}&20\cdot {\color{brown}{ 5}}&\impliedby \textit{unlimited mileag}\\ \textit{plan B}&{\color{brown}{ 5}}&{\color{blue}{ 650}}&(15\cdot {\color{brown}{ 5}})+(0.20\cdot {\color{blue}{ 650}})&\impliedby \textit{0.20 per mile} \end{array}\)

OpenStudy (mathmale):

jdoe: could you please help Tyler do his own work, instead of doing it for him? Thank you. mathmale, Moderator

OpenStudy (mathmale):

Company B: $15 per day; $0.20 per mile driven. Based upon this, how much would Company B charge for a 5 day rental, if we look at the per-day rate only? Tyler?

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

roger that didn't know.... thought he was....alrite roger that :)

OpenStudy (tylermckinney16):

$75 Mathmale.

OpenStudy (tylermckinney16):

I did $15*5

OpenStudy (mathmale):

Thank you! Tyler: Great. Yes, Company B would charge $75 for 5 days. Right. Great! Now, how much does Company B charge per mile?

OpenStudy (tylermckinney16):

And what would i do to figure that out sir?

OpenStudy (mathmale):

Sir Tyler, that info is contained in the original problem statement: "Company B: $15 per day; $0.20 per mile driven."

OpenStudy (tylermckinney16):

$0.20?

OpenStudy (mathmale):

This renter drives 650 miles, and the cost per mile is 20 cents, or $0.20. Right: $0.20 is what the renter pays per mile. But how much must he pay, TOTAL, for driving the rented car 650 miles?

OpenStudy (mathmale):

(cost per mile) * (number of miles) = total mileage charge. Multiply.

OpenStudy (tylermckinney16):

$130

OpenStudy (mathmale):

Perfect. Company B charges $75 for the 5-day rental, plus $130 for the 650 miles driven. Add these together to get the total cost payable to Company B.

OpenStudy (mathmale):

$130 + 75 ----- ?

OpenStudy (tylermckinney16):

$205

OpenStudy (mathmale):

Right. Good. How much would Company A charge? We figured this out earlier; all you have to do is to find it.

OpenStudy (mathmale):

Tyler?

OpenStudy (tylermckinney16):

Im figureing it out lol

OpenStudy (mathmale):

You've already figured it out; you just need to find the result, above.

OpenStudy (tylermckinney16):

$650?

OpenStudy (mathmale):

No, that's the number of miles driven. Look at the original problem statetment, please. How much does Company A charge per day? How much would 5 day rental cost?

OpenStudy (tylermckinney16):

$100 for 5 days? And $20 per day

OpenStudy (mathmale):

Yes, company A charges $100

OpenStudy (tylermckinney16):

So company A is the best deal.

OpenStudy (mathmale):

for 5 days. Company B charges $205. which company gets this guy's business? Yes, like you, I'd choose Company A. Done. Not too bad a problem, eh?

OpenStudy (tylermckinney16):

No u made me understand it more better.

OpenStudy (mathmale):

Delighted to hear that! Take care.